OpenStudy (anonymous):

What is the equation of the line that passes through the points (-1, 3) and (1, 11)?

OpenStudy (anonymous):

slope: (11-3) / (1-(-1) = 4 y=4x+b 1 = 4(11) + b 1 = 44 + b b = -43 equation: y= 4x-43

OpenStudy (anonymous):

But that's not in one of my options..? @jayz657

OpenStudy (anonymous):

sorry i made a mistake it is 11 = 4(1) + b 11=4 + b b = 7 equation: y= 4x +7

OpenStudy (anonymous):

The one I have similar to yours is 4x - y= -7 @jayz657

OpenStudy (anonymous):

yea its the same as the one i have